OpenStudy (anonymous):

The X-intercepts are ( -4,0) and ( 6,0 ) y-intercepts (0,240) f(x)=-10(x-p)(x-q)

OpenStudy (ranga):

When you are given the two x-intercepts,-4 and 6, you are given the two roots of f(x). So f(x) = a(x + 4)(x - 6) You are also given the y-intercept, 240 To find y-intercept, set x = 0 240 = a(4)(-6) = -24a so a = -10 So f(x) = -10(x + 4)(x - 6) is the function. If you are asked for pa dn q values, then: p = -4, q = 6 or p = 6 and q = -4.

OpenStudy (anonymous):

OOOHHH okay thank you!!

OpenStudy (ranga):

You are welcome.